For each of the following independent cases (A-E), compute the missing values in the table: Direct Manufacturing Manufacturing Total Prime CaseCost Conversion Cost Direct Materials Labor Overhead Cost 10,970 A 9,470 16,450 3,990 7,340 26,000 43,430 10,090 42,300 21,400 11,950 57,100 110,900 69,950 48,550 14,200 25,000 57,400
Answer #1

Solution :

Case Prime Cost Conversion Cost Direct materials Direct labor Manufacturing overhead Total manufacturing cost
A 9470 16450 3990 (9470-3990) = 5480 10970 (3990+5480+10970) = 20440
B (10090+7340) = 17430 (7340+26000) = 33340 10090 7340 26000 43430
C 57100 110900 42300 (57100-42300) = 14800 (110900-14800) = 96100 $153200
D (21400+11950) = 33350 48550 21400 11950 $36600 69950
E (18200+14200) = 32400 (14200+25000) = 39200 (57400-39200) = 18200 14200 25000 57400